OpenStudy (anonymous):

12 laborers could finish a roadwork in 17 days. After 11 days, 5 laborers quit and are not replaced until 4 days. (a) How many more laborers should be added to complete the job on time? (b) How much was paid for labor if during the last portion of the job, the sum of 55 was paid on wages?

OpenStudy (anonymous):

How to solve this using one unknown?

OpenStudy (shinalcantara):

since this is a work problem, you first need to find the rate of each laborer per day.. Rate will be: \[R = \frac{ 1 }{ nt }\] where R is the rate n is the number of laborers t is the time now solve for the rate given the first condition that 12 workers could finish the work in 17 days

OpenStudy (shinalcantara):

what did you get?

OpenStudy (anonymous):

So, R = 1/(12)(17)

OpenStudy (anonymous):

which is 1/204

OpenStudy (anonymous):

What should I do next?

OpenStudy (shinalcantara):

yep

OpenStudy (shinalcantara):

now we'll try to make an equation out of the next condition since it is given that 12 men were working for 11 days, 12-5 men for the next 4 days, and they need to finish the job in the next 17-11-4 days, then we'll have it as R(12men)(11days) + R(7men)(4days) + R(n)(2days) = 1 now you substitute the rate of work per worker per day

OpenStudy (shinalcantara):

1 there represents one work done

OpenStudy (anonymous):

\[\frac{40}{ 50 } +\frac{1}{ 102 }n = 1 \]

OpenStudy (anonymous):

@shinalcantara What next?

OpenStudy (shinalcantara):

try to recheck 40/50.. i think it's 40/51

OpenStudy (anonymous):

oh sorry, it is 40/51. n = 22 Is this correct?

OpenStudy (shinalcantara):

yep but since n there is the total number of workers necessary to finish the job for the last two days and there are already 7 men working at that time, therefor the addition number of workers needed will be n-7

OpenStudy (shinalcantara):

*additional

OpenStudy (anonymous):

n-7 = 22-7 = 15

OpenStudy (shinalcantara):

yep

OpenStudy (anonymous):

Ok. Thanks. What about the next question?

OpenStudy (shinalcantara):

Since you already know the number of workers for the last portion of the job, n=22, and the time is t=2days, and you're given the wage to be 55 so, payment for 1worker for 1 day would be Payment=55/(nt)

OpenStudy (shinalcantara):

The answer of that will be the rate of Payment

OpenStudy (anonymous):

Payment = 1.25

OpenStudy (shinalcantara):

To find the total payment you'll have let P be the rate of payment and Pt be the total labor cost for the entire work. so you'll have Pt = P(12)(11) + P(7)(4) + P(22)(2)

OpenStudy (anonymous):

Ok, got it! Thank you so much!

OpenStudy (shinalcantara):

ok. yw:)